# Augmenting Math Word Problems via Iterative Question Composing

### v2

@@ 3.2 MMIQC
Table 1 shows the make-up of MMIQC. MMIQC contains 3 epochs of the subsets mentioned above, except for the Mathematics Stack Exchange part.
4 Experiments
Table 2: A comparison of the accuracy of our fine-tuned models and existing models on MATH. Our evaluation result is marked with an ***. Other results, unless specified with an citation, are from (Wang et al., 2023a).
@@ 4 Experiments
Minerva-540B (Uesato et al., 2022) Arxiv+Web No maj1@64 50.3
GPT-4 Unknown No pass@1 42.4
GPT-4-0613 Unknown No pass@1 56.2
Gemini-Ultra(Team et al., 2023) Unknown No pass@1 53.2
Llama-2-7B(Touvron et al., 2023b) N/A No pass@1 2.5
Llemma-7B (Azerbayev et al., 2023) Proof-Pile-2 No pass@1 18.0
MetaMath-7B(Yu et al., 2023) MetaMathQA No pass@1 19.8
Mistral-7B-MetaMathQA(Yu et al., 2023) MetaMathQA No pass@1 28.2
@@
MAmmoTH-Coder-7B(Yue et al., 2023) MathInstruct Code pass@1 35.2
ToRA-Code-7B (Gou et al., 2023) ToRA-Corpus Code pass@1 44.6
CodeLlamma-34B N/A Code pass@1 25.0
@@
Llemma-34B-MetaMathQA MetaMathQA Shepherd maj+verify1@256 47.3
ToRA-Code-34B (Gou et al., 2023) ToRA-Corpus Code pass@1 50.8
DeepSeek-67B(Bi et al., 2024) N/A No Pass@1 18.7
@@
Deepseek-67B-MetaMathQA MetaMathQA Shepherd maj+verify1@256 48.1
4.1 Fine-tuning Setup
Our fine-tuning strategy mainly follows the practice of (Taori et al., 2023), except that we use a different prompt template to transform the question-response pairs. For a sample from Mathematics Stack Exchange, the corresponding prompt fed into the model during training is a simple concatenation of the question and response with two new-line symbols. For a sample from other subsets, we additionally add a prefix ‘Please solve the following problem and put your answer at the end with “The answer is: ”.’ to the question-response concatenation.
We use the HuggingFace transformers library (Wolf et al., 2019) for our fine-tuning experiments. We fine-tune all models on MMIQC for 1 epoch, using a $3\%$ warmup ratio linear learning rate schedule with a maximum learning rate $51e-$ 5. Employing the DeepSpeed Zero-3 Stage (Rajbhandari et al., 2020), we finetune 7B models on one node of 8xA800 gpus with micro batch-size at 8 and gradient accumulation at 4, 34B models on 2 nodes with micro batch-size at 4 and gradient accumulation at 4, maintaining an effective batch-size of 256.
4.2 Model Evaluation
We evaluate the fine-tuned models on MATH(Hendrycks et al., 2021a), a competition-level math word problems benchmark with 5000 test problems. We prompt all our fine-tuned models with the test question with the prefix ‘Please solve the following problem and put your answer at the end with “The answer is: ”.’, and extract the answer from the output using a modified version of the answer extractor provided in (Lewkowycz et al., 2022). We use a series of rules to infer whether the extracted answer is the same with the ground-truth answer, including a comparison using SymPy(Meurer et al., 2017).
The complete results of our evaluation and a comparison with existing models are shown in Table 2.
4.3 Ablation Study of Subsets of MMIQC
Table 3: How different subsets of MMIQC affect the accuracy of the finetuned model on MATH. The base model is Mistral-7B, and the accuracy baseline is the reported 28.2% of Mistral-7B-MetaMathQA.
@@ 4.3 Ablation Study of Subsets of MMIQC
+ AnsAug & QB 270.2K 30.1(+1.9)
+ AugSimilar 308.4K 31.5(+3.3)
+ IQC iter #1 330.2K 33.0(+4.8)
+ IQC iter #2 346.2K 33.7(+5.5)
+ IQC iter #3&#4 363.5K 34.4 (+6.2)
@@
In order to understand the ratio of contribution to the improvement revealed in Table 2 of different subsets of MMIQC, we fine-tune Mistral-7B with a series of training sets constructed by gradually adding the subsets. When MathStackExchange is not added, we fine-tune for 3 epochs. It can be seen from Table 3 that
@@
  • Although our filtered subset of MetaMathQA is only half the size of the original dataset (which has 395K samples, more than the total number of samples of our synthetic data), the performance drop is only 1.8%. This shows that the $k=20$ strategy in (Yu et al., 2023) results in some redundancy.
@@
  • Our Answer Augmentation & Question Boosting data help the fine-tuned model beat Mistral-7B-MetaMathQA, verifying our hypothesis that directly asking GPTs to perform question bootstrapping is more efficient than providing few-shot examples to them.
@@
  • Our IQC method leads to a significant 3.1% improvement from a high accuracy $31.5\%$ with only 55.1K samples, showing its efficientcy. Moreover, the later iterations of IQC also account for a certain ratio of improvement, proving that IQC is a method that can continuously generate new data that can help increase the diversity when added to the data generated in previous iterations.
4.4 Contamination test
@@
We check the $n$-gram matches for MMIQC to ensure that the improvement is not a result of direct memorization. We use the script provided by (Azerbayev et al., 2023) to check the $n$-gram matches between the synthetic part of the MMIQC and MATH test set. It turns out that for a 30-gram match check, there are 44 hits of match between the ‘solution’ field of MATH test set and the ‘output’ field of MMIQC, far fewer than the 168 hits between that of MATH test set and MATH training set. Moreover, we manually check these 44 hits and find that 43 among them belong to the case where intermediate steps of the solutions to similar but different questions collide, with the only exception being the question ‘A regular polygon has interior angles of 144 degrees. How many sides does the polygon have?’. This almost rules out the possibility that fine-tuned models get memorization of solutions to the problems in the test set, indicating a very low risk of data contamination for MMIQC.
5 Conclusion
@@
In this work, we introduce a novel data augmentation method for math word problem datasets called IQC and use it in the construction of our MMIQC dataset. Our evaluation results show that the models fine-tuned on MMIQC achieve new SOTAs on the MATH benchmark. The improvements of our models benefit from the diverse data source of MMIQC and the effectiveness of IQC. For future directions, we are interested in how to equip open-source models with the ability of composing questions, in order to perform IQC in a self-evolution style, similar to that in (Huang et al., 2022a). Besides, how to integrate the verification systems of (Wang et al., 2023a; Liu et al., 2023), which are originally used to improve the accuracy during inference time, into the procedure of IQC, is also an attractive topic.

### v5

@@ Introduction
Table 1 shows the make-up of MMIQC. When fine-tuning the models MMIQC contains 3 repetitions of the subsets mentioned above, except for the Mathematics Stack Exchange part. We shuffle the order of samples after combining the subsets.
Experiments
Fine-tuning Setup
Our fine-tuning strategy mainly follows the practice of (Taori et al. 2023), except that we use a different prompt template to transform the question-response pairs. For a sample from Mathematics Stack Exchange, the corresponding prompt fed into the model during training is a simple concatenation of the question and response with two new-line symbols. For a sample from other subsets, we additionally add a prefix ‘Please solve the following problem and put your answer at the end with “The answer is: ”.’ to the question-response concatenation.
We use the HuggingFace transformers library (Wolf et al. 2019) for our fine-tuning experiments.
Table 2: Ablation study on the optimal learning rate. We fine-tune Mistral-7B on MMIQC with different maximal learning rate values and evaluate the fine-tuned models on MATH to decide the best candidate.
LR 1e-6 5e-6 1e-5 2e-5 5e-5 1e-4
MATH(%) 32.3 35.1 36.0 35.4 31.5 27.1
We fine-tune all models on MMIQC for 1 epoch, using a $3\%$ warm-up ratio linear learning rate schedule. For the choice of maximum learning rate, we do a simple hyper-parameter selection experiment shown in Table 2 and determine it to be 1e-5. We use the BFloat16 numerical format during training. Employing the DeepSpeed Zero-3 Stage (Rajbhandari et al. 2020), we fine-tune 7B models on one node of 8xA800 GPUs with micro batch-size at 8, and gradient accumulation at 4, 34B models on 2 nodes with micro batch-size at 4 and gradient accumulation at 4 and $\sim$70B models on 4 nodes with micro batch-size at 4 and gradient accumulation at 2, maintaining an effective batch size of 256. It takes around 14 hours, 61 hours and 90 hours to fine-tune 7B, 34B and $\sim$70B models under the setups stated above, respectively.
Table 3: A comparative analysis of the accuracies achieved by various models on the MATH benchmark. The models marked with an asterisk(*) are fine-tuned and evaluated by us. Other results, unless otherwise cited, are derived from (Wang et al. 2023a). This comparison highlights the significant improvements our fine-tuned models demonstrate over existing solutions in mathematical problem-solving accuracy.
@@
proprietary models
   Minerva-540B (Uesato et al. 2022) Arxiv+Web No maj1@64 50.3
   GPT-4 (2023-0314) (Bubeck et al. 2023) - No pass@1 42.5
   Gemini-Ultra (Team et al. 2023) - No pass@1 53.2
$\sim$7B models
   Llama-2-7B (Touvron et al. 2023b) - No pass@1 2.5
   Qwen-7B (Bai et al. 2023) - No pass@1 11.6
   Llemma-7B (Azerbayev et al. 2023) Proof-Pile-2 No pass@1 18.0
   MetaMath-7B (Yu et al. 2023) MetaMathQA No pass@1 19.8
   Mistral-7B-MetaMathQA (Yu et al. 2023) MetaMathQA No pass@1 28.2
@@
   MAmmoTH-Coder-7B (Yue et al. 2023) MathInstruct Code pass@1 35.2
   ToRA-Code-7B (Gou et al. 2023) ToRA-Corpus Code pass@1 44.6
$\sim$34B models
   CodeLlamma-34B - Code pass@1 25.0
@@
   Llemma-34B-MetaMathQA MetaMathQA Math-Shepherd maj+verify1@256 47.3
   ToRA-Code-34B (Gou et al. 2023) ToRA-Corpus Code pass@1 50.8
$\sim$70B models
   Llama-2-70B (Touvron et al. 2023b) - No pass@1 13.5
   DeepSeek-67B (Bi et al. 2024) - No Pass@1 18.7
@@
   Deepseek-67B-MetaMathQA MetaMathQA Math-Shepherd maj+verify1@256 48.1
   Qwen-72B (Bai et al. 2023) - No pass@1 35.2
   Qwen-72B-MetaMathQA* MetaMathQA No pass@1 41.7
   Qwen-72B-MMIQC* MMIQC No pass@1 45.0
Model Evaluation
For a fair comparison, we first evaluate the fine-tuned models on MATH (Hendrycks et al. 2021a), a competition-level math word problems benchmark with 5000 test problems in a zero-shot setting. We prompt all our fine-tuned models with the test question with the prefix ‘Please solve the following problem and put your answer at the end with “The answer is: ”.’, and extract the answer from the output using a modified version of the answer extractor provided in (Lewkowycz et al. 2022). We use a series of rules to infer whether the extracted answer is the same as the ground-truth answer, including a comparison using SymPy (Meurer et al. 2017). The complete results of our evaluation on MATH and a comparison with existing models are shown in Table 3.
For the evaluation on 2023 Hungarian national high school finals in mathematics, we use the few-shot prompt used in (Paster 2023b). We manually assess the grades for every model according to the examiner instructions. The results shown in Figure 1 are the grades under a full mark of 117.
Ablation Study on Subsets of MMIQC
Table 4: How different subsets of MMIQC affect the accuracy of the finetuned model on MATH.
@@
+ AnsAug & QB +66.5K 30.1 (+1.9)
+ AugSimilar +38.2K 31.5 (+3.3)
+ IQC iter #1 +21.8K 33.0 (+4.8)
+ IQC iter #2 +16.0K 33.7 (+5.5)
+ IQC iter #3 & #4 +17.3K 34.4 (+6.2)
@@
In order to understand the ratio of contribution to the improvement revealed in Table 3 of different subsets of MMIQC, we fine-tune Mistral-7B with a series of training sets constructed by gradually adding the subsets. When MathStackExchange is not added, we fine-tune for 3 epochs. When MathStackExchange is added to the training dataset, we mix 3 repetitions of other data with 1 repetition of the MathStackExchange, and fine-tune for only 1 epoch. It can be seen from Table 4 that
@@
  • Although our filtered subset of MetaMathQA is only half the size of the original dataset (which has 395K samples, more than the total number of samples of our synthetic data), the performance drop is only 1.8%. This shows that the $k=20$ strategy in (Yu et al. 2023) results in some redundancy.
@@
  • Our Answer Augmentation & Question Boosting data help the fine-tuned model beat Mistral-7B-MetaMathQA, verifying our hypothesis that directly asking GPT to perform question bootstrapping is more efficient than providing few-shot examples to them.
@@
  • Our IQC method leads to a significant 3.1% improvement from a high accuracy of $31.5\%$ with only 55.1K samples, showing its efficiency. Moreover, the later iterations of IQC also account for a certain ratio of improvement, proving that IQC is a method that can continuously generate new data that can help increase the diversity when added to the data generated in previous iterations.
Contamination Test
@@
We check the $n$-gram matches for MMIQC to ensure that the improvement is not a result of direct memorization. We use the script provided by (Azerbayev et al. 2023) to check the $n$-gram matches between the synthetic part of the MMIQC and MATH test set. It turns out that for a 30-gram match check, there are 44 hits of match between the ‘solution’ field of MATH test set and the ‘output’ field of MMIQC, far fewer than the 168 hits between that of MATH test set and MATH training set. Moreover, we manually check these 44 hits and find that 43 among them belong to the case where intermediate steps of the solutions to similar but different questions collide, with the only exception being the question ‘A regular polygon has interior angles of 144 degrees. How many sides does the polygon have?’. This almost rules out the possibility that fine-tuned models get memorization of solutions to the problems in the test set, indicating a very low risk of data contamination for MMIQC.
Conclusion
@@
In this work, we introduce a novel data augmentation method for math word problem datasets called IQC (Iterative Question Composing) and use it in the construction of our MMIQC dataset. Our evaluation results show that the models fine-tuned on MMIQC achieve new SOTAs on the MATH benchmark. The improvements of our models benefit from the diverse data sources of MMIQC and the effectiveness of IQC.
For future directions, we are interested in how to equip open-source models with the ability to compose questions, in order to perform IQC in a self-evolution style, similar to that in (Huang et al. 2022a). Besides, how to integrate the verification systems (Wang et al. 2023a; Liu et al. 2023) that are originally used to improve the accuracy during inference time into the procedure of IQC, is also an attractive topic.
Acknowledgements
We thank Yang Yuan, Kaiyue Wen, Xingyu Dang, and Jingqin Yang for their helpful discussions.